Apply separation of variable to solve

x2uxy+9y2u=0



1
Expert's answer
2022-02-01T13:34:28-0500

Solution:

"x^2u_{xy}+9y^2u=0\ ...(i)
\\ Let\ u=f(x)g(y)
\\u_x=f'(x)g(y)
\\u_y=f(x)g'(y)"

Substituting in (i):

"x^2f'g'+9y^2fg=0
\\ x^2f'g'=-9y^2fg
\\ \dfrac{x^2f'}{f}=\dfrac{-9y^2g}{g'}=\lambda\ (\lambda>0)"

Consider "\dfrac{x^2f'}{f}=\lambda\ \& \ \dfrac{-9y^2g}{g'}=\lambda"

"\Rightarrow\dfrac{df}{f}=\dfrac{\lambda dx}{x^2}\ \&\ \dfrac{g'}{g}=\dfrac{-9y^2}{\lambda}
\\ \Rightarrow\log f=\dfrac{-\lambda}{x}+C_1\ \&\ \dfrac{dg}{g}=\dfrac{-9y^2dy}{\lambda}
\\\Rightarrow f=e^{\frac{-\lambda}{x}+C_1}\ \&\ \log g=\dfrac{-9}{\lambda}.\dfrac{y^3}{3}+C_2
\\\Rightarrow f=Ae^{\frac{-\lambda}{x}}\ \&\ \log g=\dfrac{-3y^3}{\lambda}+C_2"

"\Rightarrow g=e^{\frac{-3y^3}{\lambda}+C_2}
\\ \Rightarrow g=Be^{\frac{-3y^3}{\lambda}}
\\\therefore u(x,y)=Ae^{\frac{-\lambda}{x}}.Be^{\frac{-3y^3}{\lambda}}
\\\Rightarrow u(x,y)=Ce^{\frac{-\lambda}{x}-\frac{3y^3}{\lambda}} \ \ [C=A.B]"